MONT 106Q – Mathematical Thinking First Essay Assignment

The first of the larger paper assignments this semester will deal with the curious incident of the dog in the night-time. This essay (unlike some of the others we will do later) will deal primarily with your own personal responses to this unusual story. Your essay should show that you have done some careful reading–and some serious thinking– about the subject. You may, of course, disagree with ideas presented in class if you want. The point is for you to develop your own thinking, and then clearly express it in writing.

Your essay should be 2-3 pages in length (double-spaced, single-sided). Please submit as a MS Word .doc or .docx file to jlittle@holycross.edu by 5:00pm on Monday, September 12.

The topic and question(s) you should address:

The author of the novel, Mark Haddon, has said about his own experience writing the book that “it’s not just a book about disability ... [Christopher is] a character [who] if you met him in real life you’d never, never get inside his head. Yet something magical happens when you write a novel about him. You slip inside his head and it seems like the most natural thing in the world.” On the other hand the author Jay McInerney, in a review of the curious incident of the dog in the night-time, wrote that from his point of view as a reader of the novel, “ ... the gulf between Christopher and his parents, between Christopher and the rest of us, remains immense and mysterious ... Christopher Boone is an unsolved mystery.” Do you feel you have shared Haddon’s experience of coming to understand a truly different sort of mind? How? Are there specific incidents that helped you “get” Christopher? Or does McInerney’s sense of unsolved mystery do a better job of describing your own reactions? Is it even possible for psychologically “normal” people to understand and empathize with the experience of someone whose mind works so differently? And does this relate to our CHQ theme: If borders define us but we want to transcend them, how then shall we live?

Guidelines for this assignment

Follow these specifications:

1. It is OK to discuss preliminary ideas for what you want to say with your classmates if you want. I will also be happy to talk over things and/or read a preliminary draft to give feedback if you like.

2. To help you structure your essay, you may want to think about developing and stating a thesis in your opening paragraph – your answer(s) to the main questions above: Do you feel you have shared Haddon’s experience of coming to understand a truly different sort of mind? Or does McInerney’s sense of unsolved mystery do a better job of describing your own reactions?

3. Then, in the following paragraphs, explain in as much detail as you can why you have come to the conclusion you stated in your opening paragraph. If there are specific incidents in the novel that especially helped you understand Christopher, describe them and explain what it was that you took from them. If, on the other hand, you still feel that Christopher is totally mysterious, give some examples of his attitude or behavior that you don’t understand and explain what makes them so strange.

4. You can use quotations from the text of the curious incident of the dog in the nighttime to explain and illustrate your points. Don’t just quote them, though. Provide interpretations of those passages that come from your own thinking about them. Your evidence should be as specific, detailed, correct, and as varied as possible (focus on small passages and mention the little things that support your points).

5. When you use direct quotations, at the end place the page number(s) to identify the source. Here’s an example of what I’m looking for:

Even Christopher’s flights of imagination are curiously literal. For instance when he describes how the shape of a cloud reminds him of a spaceship, he says “... it was like an alien spaceship hundreds of kilometers long, like in Dune or Blake’s 7 or Close Encounters of the Third Kind, except that it wasn’t made of solid material, it was made of droplets of condensed water vapor, which is what clouds are made of” (curious incident, pp. 68-69).

6. You should not need to consult any other sources – this means other books, reviews of the novel, web sites, etc. This essay should present your own reactions to the novel and the main questions posed above.

7. The paper as a whole should be well-organized and “flow” from one point to the next. Do not repeat pieces of evidence unless there is a very good reason for doing so. Do not get ahead of yourself. Say one thing at a time. If you decide you want to add something, do not just tack it on somewhere. Instead, figure out where it fits in the overall scheme of your argument and put it where it belongs.

8. Your sentences should be clear, readable, and not overly wordy. You should aim for a lively style and the writing should be in your own voice. But please think twice before using very colloquial ways of saying things. If they contribute to the point you are trying to make, that is fine. But do it for a reason. Also, please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

9. Errors in mechanics (spelling, punctuation, grammar) and usage should not annoy the reader. (A word of warning: I tend to be easily annoyed by this sort of thing!) Try as much as possible to avoid those mistakes.